# Divergence-Based Similarity Function for Multi-View Contrastive Learning

### v1

@@ 5.4 More Resource Comparison
5.4.1 More Time Cost: Faster Convergence with Divergence Similarity
Self-supervised learning methods are known to require longer training times for convergence due to the nature of instance discrimination objectives and the inherent challenge of learning meaningful representations from unlabeled data. This issue becomes even more pronounced when dealing with complex datasets or limited batch sizes.
To investigate convergence behavior, we extended the number of training epochs and compared model accuracy as a function of elapsed training time, as shown in Figure 3. While the baseline MoCo gradually improves over a long training schedule (55 hours at 8000 epochs), our proposed DSF reaches comparable accuracy within only 11 hours at 400 epochs, using merely 20% of the total training time. Specifically, compared to the Loss_avg and Feature_avg baselines trained for 800 epochs, our DSF reaches similar accuracy within just 400 epochs, achieving a 2× speedup in convergence. Moreover, we observe that the performance gap between our DSF and the two other multi-view baselines (Loss_avg, Feature_avg) remains consistently large, regardless of the total training time. This indicates that our method not only converges faster but also achieves a higher final accuracy that is not easily matched even with extended training.
This result highlights that our divergence-based formulation facilitates more efficient optimization in self-supervised learning. Notably, all three multi-view methods considered in this comparison have identical per-epoch training time, ensuring a fair evaluation.
5.4.2 More GPU Memory : Number of Views
We sample multiple augmented views from each image; thus, it is important to investigate how the number of sampled views affects the quality of the learned representations. Prior studies [Tian et al., 2020, Shidani et al., 2024, Pototzky et al., 2022] have shown that increasing the number of augmentations generally leads to better representation quality.
In our framework, the number of views directly impacts the accuracy of the estimated von Mises–Fisher (vMF) distribution. As the number of views increases, the estimated parameters $\mu$ and $\kappa$ become more reliable due to reduced estimation variance.
Figure 4 presents the performance trends of three multi-view learning methods on both kNN and linear evaluation tasks, as the number of views $M$ increases. Across all view counts, DSF consistently outperforms the other two approaches. These results suggest that our divergence-based formulation provides stable and reliable improvements in multi-view learning settings.
@@ 5.5 Temperature Free
In numerous studies, the temperature parameter $\tau$ has been shown to be critical for achieving optimal performance in contrastive learning [Zhang et al., 2021, Zhang et al., 2022, Kukleva et al., 2023]. In our method, however, $\tau$ is not required, as discussed in Proposition 1, and it can even degrade performance when introduced.
To validate this, we conducted experiments over a wide range of temperature values, from 0.1 to 5.0. Interestingly, setting $\tau=1$ (i.e., using the raw similarity without scaling) consistently resulted in the best performance in Table 3. This suggests that our DSF is inherently numerically stable and does not require additional temperature tuning.
Temperature $\tau$ 0.8 0.9 1 1.1
kNN 91.2 91.3 91.8 91.4
Linear Evaluation 92.0 92.3 92.5 92.0
Table 3: Temperature Experiment. MoCo + DSF results in CIFAR10.
5.6 Training Details
To train on ImageNet-100 dataset, MoCo v3 with ViT model as the encoder from scratch. We use AdamW as our optimizer. The weight decay is 0.1 and the encoder momentum is 0.99 for updating encoder. We choose batch size of 512, learning rate of 0.03. In our proposed method, a larger learning rate of 0.06 is used. GPU used for overall training is NVIDIA GeForce RTX 3090.
6 Future Work
@@
We use von Mises-Fisher distribution and Kullback-Leibler divergence for our proposed divergence-based similarity function. Although they are popular choice, there are plenty of available distributions and divergences. For example, in divergence, there are other divergences such as Rényi divergence, $\chi^{\alpha}$divergence, and Jensen–Shannon divergence—collectively known as the family of $f$-divergences could be explored. Investigating how these alternative divergences influence the representation quality and computational efficiency of contrastive methods remains a promising direction for future work.
@@ Acknowledgements
This work is supported by the National Research Foundation of Korea(NRF) grant funded by the Korea government(MSIT) (RS-2024-00421203, RS-2024-00406127) and MSIT/IITP(2021-0-00077)
@@ Appendix B Feature Averaging Also Encodes Only Pairwise Relationships
Appendix C Stabilization of the Concentration Parameter $\kappa$
To address numerical instability during training, we use the approximated concentration parameter $\kappa$ as defined in Equation 5. However, even this approximation can lead to excessively large values of $\kappa$ as training progresses. To mitigate this, we propose two stabilization techniques.
Figure 6 illustrates the original approximated $\kappa$ values computed from Equation 5, alongside the stabilized values obtained after applying our proposed techniques.
Figure 6: Consistency by Kappa Normalization. $\kappa$ value is calculated as approximated formulation as (5). $x$-axis is $\bar{R}$ and $y$-axis is calculated $\kappa$. $p$ is feature dimension size. In our backbone ResNet18, feature dimension is 128. (upper) Directly calculated $\kappa$. (lower) $\kappa$ is divided by feature dimension $p$ and calculated with $\bar{R}$ multiplied by $\lambda_{\bar{R}}$. Without these two techniques, $\kappa$ value is going excessively large.

### v2

@@ 5.4 More Resource Comparison
More Time Cost: Faster Convergence with Divergence Similarity
Self-supervised learning often requires longer training to converge, due to the difficulty of learning meaningful representations without labels (Chen et al. 2020a; Wang et al. 2021a, b). To investigate convergence behavior, we compare model accuracies over elapsed training time (Figure 3). While MoCo requires 55 hours for 8000 epochs, DSF achieves similar accuracy in just 11 hours (400 epochs), using only 20% of the time. Compared to Loss_avg and Feature_avg, DSF also reaches similar accuracy with half the epochs and maintains a consistently higher performance throughout training. These results highlight the efficiency of DSF in optimizing self-supervised objectives.
More GPU Memory: Consistent Superiority Across View Counts
In Figure 4, we examine how the number of augmented views affects representation quality, as increasing views generally improves performance (Tian, Krishnan, and Isola 2020; Shidani et al. 2024; Pototzky, Sultan, and Schmidt-Thieme 2022). In DSF, more views lead to more accurate estimation of the vMF parameters $\mu$ and $\kappa$, due to reduced variance. As shown in Figure 4, DSF consistently outperforms other multi-view methods across different numbers of views, clearly demonstrating the superiority of DSF in multi-view representation learning.
@@ 5.5 Temperature Free
While prior works (Zhang et al. 2021, 2022; Kukleva et al. 2023) emphasize the importance of the temperature parameter $\tau$ in CL, DSF does not require it (see Proposition 1) and even shows degraded performance when $\tau$ is introduced. As shown in Table 4, the best results are consistently achieved when $\tau=1$, indicating that DSF is inherently stable without temperature tuning.
$\tau$ 0.1 0.9 1 1.1 5
kNN 27.71 89.83 90.04 89.80 38.44
Lin. Eval. 22.02 91.18 91.21 91.09 56.33
Table 4: Temperature Experiment. Training MoCo with DSF under varying temperatures on CIFAR-10.
@@
6 Conclusion
In this work, we propose a novel similarity function, named the Divergence-based Similarity Function (DSF), which captures joint relationships among multiple features. Our method demonstrates strong performance under both equal and increased resource settings, surpassing baseline and other multi-view methods. In addition, it produces highly transferable and robust representations, as evidenced by its effectiveness in transfer learning and resilience to data corruptions. We also establish its connection to cosine similarity and demonstrate that DSF performs best with a fixed temperature of 1, eliminating the need for tuning.
DSF is defined using the von Mises–Fisher distribution and Kullback–Leibler divergence. While effective, other $f$-divergences such as Rényi, $\chi^{\alpha}$, and Jensen–Shannon divergence remain unexplored, offering promising directions for future research.
Beyond these avenues, a natural extension is a deeper theoretical study of DSF. In particular, examining its generalization behavior and convergence in high-dimensional regimes could provide stronger foundations for its empirical success. Such analysis may also clarify why DSF exhibits superior robustness and transferability compared to cosine similarity.
@@ Acknowledgements
This work is supported by the National Research Foundation of Korea(NRF) grant funded by the Korea government(MSIT) (RS-2024-00421203, RS-2024-00406127, NRF-2016K2A9A2A13003815) and IITP(2021-0-00077). We would also like to thank Imseong Park, Hyunsung Bae, Jongyeon Lee and Jaesuk Kim, for their valuable feedback on an earlier version of this manuscript.
@@ Appendix B Feature Averaging Also Encodes Only Pairwise Relationships
Appendix C Further Explanation of DSF
C.1 Definitions
In here, we introduce the definition of $A_{p}(\cdot)$ and $I_{p}(\cdot)$ in Eq. (3) and Eq. (LABEL:eq:kld) respectively. Here, $\Gamma$ denotes the gamma function. $I_{p}(\cdot)$ is Modified Bessel function of the first kind with order $p$. $A_{p}(\cdot)$ is a fraction of two Modified Bessel function of the first kind.
$\begin{split}&I_{p}(\kappa)=\sum_{m=0}^{\infty}\frac{1}{m!\Gamma(m+p+1)}(\frac{\kappa}{2})^{2m+p},\\ &A_{p}(\kappa)=\frac{I_{\frac{p}{2}}(\kappa)}{I_{\frac{p}{2}-1}(\kappa)}\end{split}$ (20)
C.2 Second Term Calculation
In Eq. (LABEL:eq:kld), the second term cannot be evaluated in closed form using the modified Bessel function $I_{p}(\cdot)$. Therefore, we adopt an asymptotic approximation as given in (DLMF , Eq. (10.41.3)).
[…]
Using this approximation, we compute the second term as follows, truncating the series at $k=5$.
$\begin{split}&\log\frac{I_{\frac{p}{2}-1}(\kappa_{1})}{I_{\frac{p}{2}-1}(\kappa_{0})}=\log\frac{I_{v}(vz_{1})}{I_{v}(vz_{0})}\\ &=v(\eta_{1}-\eta_{0})+\frac{1}{4}(\frac{1+z_{0}^{2}}{1+z_{1}^{2}})+\log\sum_{k=0}^{\infty}\frac{U_{k}(x_{1})}{v^{k}}\cdot\frac{1}{{\sum_{k=0}^{\infty}\frac{U_{k}(x_{0})}{v^{k}}}}\\ &v=\frac{p}{2}-1\end{split}$ (22)
C.3 Stabilization of the Concentration Parameter $\kappa$
To address numerical instability during training, we use the approximated concentration parameter $\kappa$ as defined in Eq. (5). However, even this approximation can lead to excessively large values of $\kappa$ as training progresses. To mitigate this, we propose two stabilization techniques.
Figure 6 illustrates the original approximated $\kappa$ values computed from Eq. (5), alongside the stabilized values obtained after applying our proposed techniques.
Figure 6: Consistency by Kappa Normalization. $\kappa$ value is calculated as approximated formulation as Eq. (5). $x$-axis is $\bar{R}$ and $y$-axis is calculated $\kappa$. $p$ is feature dimension size. In our backbone ResNet18, feature dimension is 128. (upper) Directly calculated $\kappa$. (lower) $\kappa$ is divided by feature dimension $p$ and calculated with $\bar{R}$ multiplied by $\lambda_{\bar{R}}$. Without these two techniques, $\kappa$ value is going excessively large.
Appendix D Experimental Details
D.1 Training Details
To train on ImageNet-100 dataset, MoCo v3 with ViT model as the encoder from scratch. We use AdamW as our optimizer. The weight decay is 0.1 and the encoder momentum is 0.99 for updating encoder. We choose batch size of 512, learning rate of 0.03. In DSF, a larger learning rate of 0.06 is used. GPU used for overall training is NVIDIA GeForce RTX 3090.
D.2 Transfer Learning
We conduct transfer learning experiments following the protocol in (Chen et al. 2020a). We report top-1 accuracy for CIFAR-10, CIFAR-100 and the Describable Textures Dataset (DTD); and mean per-class accuracy for FGVC Aircraft and Oxford-IIIT Pets.
D.3 Robustness under Distribution Shift
We adopt the corruption methodology introduced by  (Hendrycks and Dietterich 2019). The ImageNet-100-C dataset is constructed by extracting the 100 selected classes from the original ImageNet-C benchmark. We evaluate the models using the mean Corruption Error (mCE) metric on corrupted images with severity level 1.
The evaluation covers all 18 corruption types provided in the benchmark: gaussian noise, shot noise, impulse noise, defocus blur, glass blur, motion blur, zoom blur, snow, frost, brightness, contrast, elastic transform, pixelate, jpeg compression, speckle noise, gaussian blur, spatter, and saturate.
(a) Three terms and KL divergence value of positive pairs.
(b) Three terms and KL divergence value of negative pairs.
Figure 7: (Top) Positive pairs. (Bottom) Negative pairs. Positive pairs maintain KL divergence values close to 0, reflecting their similarity, while negative pairs exhibit larger divergence values, reflecting increasing dissimilarity.